1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Christian Morgenstern,

The Banshee (An Approach)

La lupfanto- mo El Hechicero

The Werewolf

translated by Max Knight translated by Bertilo Wennergren translated by Hans Eiseneck translated by Alexander Gross

One night, a banshee

slunk away from mate and child,

and in the gloom went to a village tea-

cher’s tomb, requesting him: “Inflect

me, pray.”

Jen lupfantom’ de

famili’ meznokte iam ku-

ris for al tombo de la pro-

fesor’. kaj petis pri konju-

gaci’.

Un hechicero, con

esposa vino un día a la fo-

sa de su maestro, y al

profesor pidió: “¿Me cuen-

tas? ¡Por favor!”

A Werewolf, troubled by his

name, Left wife and brood one

night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

The village teacher

climbed up straight upon his grave stone

with its plate and to the apparition

said who meekly knelt before

the dead:

Do diktis tiu dokta

ul’ starante sur la

kructabul’ dum lup’ senmova

kiel splinto atentis vortojn de

l’ mortinto.

El profesor sin va-

cilar subió y comenzó a

hablar. Desde allá de su le-

trero se dirigió al hechi-

cero:

“Oh sage, wake up, please

don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed on-

ce, and then was ready.

“The banSHEE, in the

subject’s place; the banHERS, the pos-

sessive case. The banHER, next, is

what they call objective case–and that

is all.”

De “lupfantomo”,

simpla rekta, tra “lupfintomo”,

plej perfekta, plu “lupfontomo”,

iom rara, ĝis “lupfuntomo”,

ho, koŝmara.

“El hechiuno” le

explica, “el hechidos” de-

spués indica, “el hechitres” dice

al fin, “con eso basta ¡chi-

quitín!”

“Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singular-

ly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in

this same sense.”

The banshee marveled

at the cases and writhed with plea-

sure, making faces, but said: “You did not

add, so far, the plural to the singu-

lar!”

Ĉi participoj plaĉis

multe al la fantomo, kaj

rezulte ĝi petis: “Legu plu

el PIV-o pri participoj en

pasivo.”

El hechicero, con-

movido por lo que capta su

oído, al maestro ruega,

encantado, que le adjunte un

quebrado.

“I know that–I’m no mental

cripple– The future form and parti-

ciple Are what I crave,” the beast

replied. The scholar paused–again

he tried:

…  …  …  …

2/3

The teacher, though,

admitted then that this was not within

his ken. “While bans are fre-

quent”, he advised, “A she cannot be pluri-

zed.”

La viro ĉagreniĝis

tamen: “Jam fermis mi per

fin’ kaj amen, ĉar netransiraj be-

stoj feblas kaj la pasiv’ ne-

niam eblas.”

Pero el profesor

confiesa que nada sabe con

certeza: “Hay muchos he-

chos. Hechiceros

únicamente hay enteros.”

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no

future.”

The banshee, rising

clammily, wailed: “What about my

family?” Then, being not a lear-

ned creature, said humbly “Thanks”

and left the teacher.

Sed lupfantomo pri

transiro ja spertas pli ol eĉ

vampiro. La lup’ revenis lar-

mokula al sia famili’ ulula.

El brujo se volvió

muy triste: “Mi cara mitad,

¿no la viste?” Ya que no era sa-

bio con gratitud se de-

spidió.

The Werewolf knew better– his sons still slept At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

…  …  …  …

3/3

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Esperanto by Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu la retejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into hispana by Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .